What volume is occupied by 0.897 mol at 394.9 K and 861 mmHg? ___L
Answer #1

Answer:-

n= 0.897 mol

T = 394.9 K

P = 861 mmHg

= 861/760 atm ( 1 atm = 760 mmHg)

PV = nRT

V = nRT/P

= 0.897 x 0.0821 x 394.9 /(861/760)

= 0.897 x 0.0821 x 394.9 x 760/861

=25.67 L